# Non-Euclidean Erdős–Anning Theorems

### v2

@@ 1 Introduction
A well-known theorem of Paul Erdős and Norman Anning states that every point set in the Euclidean plane with integer distances must be either collinear or finite [1, 11]. More strongly, for any (non-degenerate) triangle of diameter $\delta$, at most $O(\delta^{2})$ points can have integer distances from all three triangle vertices. When the whole point set has diameter $D$, its number of points is $O(D)$ [29]. In contrast, as Euler already observed, there exist infinite non-collinear point sets of bounded diameter with all distances rational, for instance, dense subsets of a unit circle [12, 14].
There are many metrics other than Euclidean distance for which we might ask similar questions. One obstacle to generalizing the Erdős–Anning theorem to other metrics is that its usual proof is heavily based on algebraic geometry. In the Euclidean plane, the locus of points whose distances to two given points differ by a fixed integer is a branch of a hyperbola. Therefore, the points with integer distances to three non-collinear points lie at the intersection points of two finite families of hyperbolas. Hyperbolas are algebraic curves of degree two, and by Bezout’s theorem two hyperbolas have at most four crossings, so these two finite families have finitely many intersection points. But it is not obvious how to extend this argument to metrics for which the the corresponding curves are more complicated or non-algebraic. Erdős and Anning originally used a different argument using trigonometric inequalities [1], but its generalization is also non-obvious.
A second obstacle is, of course, that analogous statements are untrue for some metrics. For the $L_{1}$ or $L_{\infty}$ distances in the plane, for instance, the integer lattice provides a familiar example of an infinite set that is not collinear and has only integer distances. Thus, generalizations of the theorem to other families of metrics must avoid families that include $L_{1}$ or $L_{\infty}$.
@@
For any two-dimensional complete Riemannian manifold of bounded genus $g$, every point set with integer distances must be either be finite or have shortest curves through each triple of points that all lie on a single geodesic. If some three of the points do not lie on a single shortest curve, and have diameter $\delta$, at most $O((g+1)\delta^{2})$ points can have integer distances from these three points. (Section 3)
For geodesic distance on the boundary of a convex set in $\mathbb{R}^{3}$, every point set with integer distances must either be finite or have shortest curves through each triple of points that all lie on a single geodesic. If some three of the points do not lie on a single shortest curve, and have diameter $\delta$, at most $O(\delta^{2})$ points can have integer distances from these three points. If a set with integer distances has diameter $D$, it has at most $O(D^{4/3})$ points. (Section 4)
Our proofs replace the intersection properties of low-degree algebraic curves with the geometric and topological properties of additively weighted Voronoi diagrams. Our results for Riemannian manifolds resolve a question posed in 1983 by Richard Guy [13], on the equilateral dimension of Riemannian manifolds, the maximum number of points at unit distance from each other in these manifolds. For complete Riemannian 2-manifolds of bounded genus, we bound the equilateral dimension by a function of the genus (Corollary 20) but for locally Euclidean incomplete Riemannian 2-manifolds, for complete Riemannian 2-manifolds of unbounded genus, and for complete Riemannian metrics on $\mathbb{R}^{3}$, the equilateral dimension is unbounded (Example 6 and Example 7).
@@ 2 Convex distance functions
when this limit exists. Convex distance functions define geodesic metric spaces, spaces in which the distance between any two points equals the length of a shortest curve between the points. For strictly convex distance functions, this shortest curve is unique: it is just the line segment between its two points. It is convenient to combine these observations about shortest curves with the triangle inequality to obtain the following observation about distances:
@@ 2.1 Examples
2.2 Voronoi diagrams and star-shaped cells
A Voronoi diagram of a finite set of points (called sites) in a metric space $(X,d)$ subdivides the space into subsets according to which of the sites is closest. More precisely, each site $s_{i}$ has a (closed) cell $V_{i}$ associated with it, where
is the set of points in the metric space for which the distance to site $s_{i}$ is less than or equal to the distance to any other site $s_{j}$. We will only use Voronoi diagrams of finite sets (in fact, of at most three sites), so we only define and discuss the properties of these diagrams for finite sets of sites.
Voronoi diagrams have been the object of intensive study, including Voronoi diagrams for convex distance functions [7, 16, 20], for points on a sphere [24, 3, 26], for points in the hyperbolic plane [9, 28, 4], orbifolds [23], and for more general Riemannian manifolds [21]. Additively weighted Voronoi diagrams are also a standard concept [6, 17], and confusingly, are sometimes called hyperbolic Dirichlet tessellations despite their use of Euclidean distance [2], because the boundaries of their cells are arcs of hyperbolas (the same hyperbolas used in the Erdős–Anning proof). We did not find works combining both additive weights and non-Euclidean distances.
@@ 2.2 Voronoi diagrams and star-shaped cells
2.3 Degenerate and non-degenerate sites
In unweighted Voronoi diagrams, every cell is a topological disk (or the whole space), having its site in its interior. That is not true for additively weighted Voronoi diagrams, but the cells that do not have their site in their interior have a very special structure, which we examine next.
@@ 2.4 Intersections of non-degenerate cells
In the previous subsection, we studied the shapes of individual cells of additively weighted Voronoi diagrams. In this section we instead consider the intersection patterns of cells.
@@
Suppose for a contradiction that there existed a point $p$ interior to $V_{i}$ that also belonged to $V_{j}$. Then segments $ps_{i}$ and $ps_{j}$ must be disjoint except for their shared endpoint, for otherwise one would be a subset of the other, and one of the two sites would belong to the Voronoi cell of the other, contradicting the assumption that both are non-degenerate. Because $p$ belongs to both $V_{i}$ and $V_{j}$, it has equal weighted distance to both sites. Because $p$ is interior to $V_{i}$, we can extend line segment $ps_{i}$ within a neighborhood of $p$ inside $V_{i}$, to obtain a line segment $qs_{i}$ that still remains entirely interior to $V_{i}$ (Fig. 7). By the Lipschitz property of distances, $q$ is farther than $p$ from $s_{i}$ by an amount equal to the distance from $p$ to $q$, while the distance from $q$ to $s_{j}$ cannot increase as quickly. Therefore, $q$ belongs to the Voronoi diagram of $V_{j}$ but not of $V_{i}$, contradicting our construction of $q$ as a point in $V_{i}$. This contradiction shows that our assumption of a point interior to $V_{i}$ that also belonged to $V_{j}$ cannot hold: no such point exists. Symmetrically, no point interior to $V_{j}$ can also belong to $V_{i}$. ∎
@@ 2.5 Erdős–Anning theorems for strict convex distance functions
Weight $w_{1}$ is identically zero, and by the triangle inequality the remaining two weights have absolute values $|d(d_{1},p)-d(s_{2},p)|$ and $|d(s_{1},p)-d(s_{3},p)|$ that are at most $d(s_{1},s_{2})$ and $d(s_{1},s_{3})$ respectively. Therefore, there are at most $\bigl{(}2d(s_{1},s_{2})+1\bigr{)}\cdot\bigl{(}2(d(s_{1},s_{3})+1\bigr{)}$ combinations of weights defining an additively weighted Voronoi diagram for which $p$ might be a triple intersection point, and at most two triple intersection points per diagram, giving a total of at most $2\bigl{(}2d(s_{1},s_{2})+1\bigr{)}\cdot\bigl{(}2(d(s_{1},s_{3})+1\bigr{)}$ possible points that can have integer distances to all three of $s_{1}$, $s_{2}$, and $s_{3}$. ∎
As an immediate corollary, we have the following generalization of the Erdős–Anning theorem:
@@
If $S$ is not collinear, it has three non-collinear points to which we can apply Theorem 1. ∎
@@ 2.6 Diameter bounds
Let $\rho$ be the minimum Euclidean distance to the origin of the unit ball of the given distance function $d$. Then scaling both $S$ and the distance function by $1/\rho$ produces a distance function whose unit ball contains a given circle, without changing the distance between any two of the scaled points in $S$. Let $\xi$ be the maximum Euclidean distance from the origin of this scaled unit ball; note that, if we treat the distance function as fixed, then $\xi$ is a constant. After this scaling, the function measured by $d$ is upper bounded by the Euclidean distance, and lower bounded by Euclidean distance divided by $\xi$. Therefore, the Euclidean diameter of $S$ (as scaled) is also at most $D$.
@@
Enclose $S$ in a square bounding box of (Euclidean) side length $D$, and center this within a larger square $B$ of side length $3D$, so that all points of $S$ are at Euclidean distance at least $D$ from the boundary of $B$. Construct the (Euclidean) Voronoi diagram of $S$, let $s_{1}$ be the point of $S$ whose Voronoi cell $V_{1}$, intersected with $B$, has the smallest Euclidean area, $\leq 9D^{2}/n=O(D^{2}/n)$. Let $s_{2}$ be the Euclidean nearest neighbor of $s_{1}$ in $S$, let $m$ be the midpoint of $s_{1}$ and $s_{2}$, and construct a right triangle $ps_{1}m$ with right angle at $s_{1}$, contained in $V_{1}\cap B$, with the largest possible Euclidean area. (See Fig. 9 for an example.) We distinguish two cases:
  • If $p$ is on the boundary of $B$, as shown in Fig. 9, it is at Euclidean distance $\geq D$ from $s_{1}$. In order for right triangle $ps_{1}s_{2}$ to have area $O(D^{2}/n)$, at most the area of the set $V_{1}\cap B$ that contains it, edge $s_{1}s_{2}$ must have Euclidean length $O(D/n)$, and therefore also $d(s_{1},s_{2})=O(D/n)$. For this distance to be a nonzero integer, $n$ must be $O(D)$.
  • Otherwise, $p$ is on the boundary of $V_{1}$, and equidistant (in Euclidean distance) from $s_{1}$ and some site $s_{3}$ which cannot be collinear with $s_{1}s_{2}$. Because the Euclidean distances and distances according to $d$ are proportional, $d(s_{1},s_{2})\cdot d(s_{1},s_{3})=O(D^{2}/n)$. It follows from Theorem 1 that $n=O(D^{2}/n)$ and therefore that $n=O(D)$.∎
@@ 3 Riemannian manifolds
One complication in these spaces is that their usual analogues of straight lines, geodesics, are only locally shortest curves between their points: two or points may belong to a common geodesic that does not contain the shortest curve between them. Indeed, in some cases a geodesic may cross itself or form a dense subset of its surface. For this reason, we will generally work with shortest curves (between some two points of the space) rather than geodesics. These curves may not be unique: for instance, for antipodal points on a sphere, there are infinitely many shortest curves, covering the whole sphere. By the Hopf–Rinow theorem [15], every complete Riemannian manifold has a shortest curve between every two points. Thus, these are geodesic metric spaces, and in particular the Lipschitz property of distances (1) applies to them.
@@ 3.1 Examples
For any $n$, one may form a complete Riemannian manifold from $n$ spheres, connected by cutting out disks from pairs of spheres, attaching a cylindrical tube between the boundaries of the two disks, and smoothing the parts of the surface where the disks meet the tubes. This construction may be performed abstractly as a topological space; it is not necessary to embed the spheres and tubes into three-dimensional space, allowing the lengths of the tubes to be adjusted. Let $S$ consist of one point per sphere, for spheres small enough that all distances within one sphere are less than $1/3$; then by adjusting the lengths of all the tubes, all distances between points of $S$ can be made equal to one. This construction produces arbitrarily large finite sets of points at unit distance from each other, with genus $O(n^{2})$. Because they all have the same distance, no three of these points belong to a shortest curve.
The connected sum of an infinite sequence of examples of this type for $n=4,5,6,\dots$, obtained by using tubes to connect the example for each $n$ to the examples for $n-1$ and $n+1$, forms a single complete Riemannian manifold of infinite genus in which one can find arbitrarily large (but nevertheless finite) sets of points that are all at unit distance from each other.
@@
Our proof of the Erdős–Anning theorem for strictly convex distance functions relied on the non-planarity of the graph $K_{3,3}$. For complete Riemannian manifolds of bounded genus we will need to reach farther for graphs that cannot be embedded, as the following example shows.
@@
Form a complete Riemannian manifold (the hexagonal torus) by gluing opposite edges of a Euclidean regular hexagon. After this gluing step, the six vertices of the hexagon become two points of the hexagonal torus. Let $S$ be the set consisting of these two points together with the center of the hexagon. Then the hexagonal torus has six points equidistant from the three points of $S$. Connecting the three points of $S$ by shortest curves to each of these six points forms an embedding of the graph $K_{3,6}$ on the hexagonal torus (Fig. 11).

### v3

@@ 1 Introduction
Many incidence properties of lines or curves, seemingly mysterious on their own, become obvious when viewed through the lens of computational geometry, where the curves may be features of Voronoi diagrams or related constructions. Thus, for instance, to see that the three perpendicular bisectors of a triangle’s sides meet at a single point, observe that these bisectors contain the edges of the Voronoi diagram of the triangle vertices, and these Voronoi edges meet at the Voronoi vertex. In this work we apply the same principle to the Erdős–Anning theorem on integer distances in the plane, conventionally proven using the algebraic intersection properties of hyperbolae. By reinterpreting these hyperbolae and their intersection points as the cell boundaries and vertices of additively-weighted Voronoi diagrams, and by analyzing these diagrams geometrically and topologically rather than algebraically, we greatly generalize this theorem, to several wide classes of two-dimensional metric spaces. In doing so, we resolve a 1983 question of Richard Guy on the equilateral dimension of Riemannian manifolds, the maximum number of points that can be placed at equal distances from each other in these spaces.
The Erdős–Anning theorem states that every point set in the Euclidean plane with integer distances must be either collinear or finite [1, 11]. More strongly, for any (non-degenerate) triangle of diameter $\delta$, at most $O(\delta^{2})$ points can have integer distances from all three triangle vertices. When the whole point set has diameter $D$, its number of points is $O(D)$ [30], and if in addition it is non-collinear its number of points is $D^{O(1/\log\log D)}$ [13]. In contrast, as Euler already observed, there exist infinite non-collinear point sets of bounded diameter with all distances rational, for instance, dense subsets of a unit circle [12, 15].
There are many metrics other than Euclidean distance for which we might ask similar questions. One obstacle to generalizing the Erdős–Anning theorem to other metrics is that its usual proof is heavily based on algebraic geometry. In the Euclidean plane, the locus of points whose distances to two given points differ by a fixed integer is a branch of a hyperbola. Therefore, the points with integer distances to three non-collinear points lie at the intersection points of two finite families of hyperbolas. Hyperbolas are algebraic curves of degree two, and by Bezout’s theorem two hyperbolas have at most four crossings, so these two finite families have finitely many intersection points. But it is not obvious how to extend this argument to metrics for which the corresponding curves are more complicated or non-algebraic. Erdős and Anning originally used a different argument using trigonometric inequalities [1], but its generalization is also non-obvious.
A second obstacle is that analogous statements are untrue for some metrics. For the $L_{1}$ or $L_{\infty}$ distances in the plane, the integer lattice provides a familiar example of an infinite set that is not collinear and has only integer distances. Thus, generalizations of the theorem to other families of metrics must avoid families that include $L_{1}$ or $L_{\infty}$.
@@
For any two-dimensional complete Riemannian manifold of finite genus $g$, every point set with integer distances must either be finite or be contained in a geodesic (a locally-shortest curve) that contains every shortest curve between pairs of the points. If some three points do not lie on a single shortest curve, and have diameter $\delta$, at most $O((g+1)\delta^{2})$ points can have integer distances from these three points. (Section 4)
For geodesic distance on the boundary of a convex set in $\mathbb{R}^{3}$, every point set with integer distances must either be finite or be contained in a geodesic that contains every shortest curve between pairs of the points. If some three points do not lie on a single shortest curve, and have diameter $\delta$, at most $O(\delta^{2})$ points can have integer distances from these three points. If a set with integer distances has diameter $D$, it has at most $O(D^{4/3})$ points. (Section 5)
Our proofs replace the intersection properties of low-degree algebraic curves with the geometric and topological properties of additively weighted Voronoi diagrams. Our results for Riemannian manifolds resolve a question posed in 1983 by Richard Guy [14], on the equilateral dimension of Riemannian manifolds, the maximum number of points at unit distance from each other in these manifolds. Guy asked whether the equilateral dimension could be bounded by a function of dimension. For complete Riemannian 2-manifolds of bounded genus, we bound the equilateral dimension by a function of the genus (Corollary 20) but for locally Euclidean incomplete Riemannian 2-manifolds, for complete Riemannian 2-manifolds of unbounded genus, and for complete Riemannian metrics on $\mathbb{R}^{3}$, we show that the equilateral dimension is unbounded (Example 6 and Example 7).
2 Preliminaries: Voronoi diagrams and star-shaped cells
A Voronoi diagram of a finite set of point sites in a metric space $(X,d)$ subdivides the space according to which site is closest. Each site $s_{i}$ has a (closed) cell $V_{i}$ associated with it, where
is the set of points whose distance to site $s_{i}$ is less than or equal to the distance to any other site $s_{j}$. We will only use Voronoi diagrams of finite sets (in fact, of at most three sites), so we only define and discuss the properties of these diagrams for finite sets of sites.
Voronoi diagrams have been the object of intensive study, including Voronoi diagrams for convex distance functions [7, 17, 21], for points on a sphere [25, 3, 27], for points in the hyperbolic plane [9, 29, 4], orbifolds [24], and for more general Riemannian manifolds [22]. Additively weighted Voronoi diagrams are also a standard concept [6, 18], and confusingly, are sometimes called hyperbolic Dirichlet tessellations despite their use of Euclidean distance [2], because their cells are bounded by hyperbolas (the same hyperbolas from the Erdős–Anning proof). We did not find works combining additive weights and non-Euclidean distances.
@@ 3 Convex distance functions
when this limit exists. Convex distance functions define geodesic metric spaces in which the distance between any two points equals the length of a shortest curve between the points. For strictly convex distance functions, this shortest curve is unique: it is just the line segment between its two points. It is convenient to combine these observations about shortest curves with the triangle inequality to obtain the following observation about distances:
@@ 3.1 Examples
3.2 Degenerate and non-degenerate sites
In unweighted Voronoi diagrams, every cell is a topological disk (or the whole space), having its site in its interior. That is not true for additively weighted Voronoi diagrams, but the cells that do not have their site in their interior have a very special structure, which we examine in this section.
@@ 3.3 Intersections of non-degenerate cells
In Section 3.2, we studied the shapes of individual cells of additively weighted Voronoi diagrams. In this section we instead consider the intersection patterns of cells.
@@
Suppose for a contradiction that there existed a point $p$ interior to $V_{i}$ that also belonged to $V_{j}$. Then segments $ps_{i}$ and $ps_{j}$ must be disjoint except for their shared endpoint, for otherwise one would be a subset of the other, and one of the two sites would belong to the Voronoi cell of the other, contradicting the assumption that both are non-degenerate. Because $p$ belongs to both $V_{i}$ and $V_{j}$, it has equal weighted distance to both sites. Because $p$ is interior to $V_{i}$, we can extend line segment $ps_{i}$ within a neighborhood of $p$ inside $V_{i}$, to obtain a line segment $qs_{i}$ that still remains entirely interior to $V_{i}$ (Fig. 7). By the Lipschitz property of distances, $q$ is farther than $p$ from $s_{i}$ by an amount equal to the distance from $p$ to $q$, while the distance from $q$ to $s_{j}$ cannot increase as quickly. Therefore, $q$ belongs to the Voronoi diagram of $V_{j}$ but not of $V_{i}$, contradicting our construction of $q$ as a point in $V_{i}$. This contradiction shows that our assumption of a point interior to $V_{i}$ and in $V_{j}$ cannot hold: no such point exists. Symmetrically, no point interior to $V_{j}$ can also belong to $V_{i}$. ∎
@@ 3.4 Erdős–Anning theorems for strict convex distance functions
Weight $w_{1}$ is identically zero, and by the triangle inequality the remaining two weights have absolute values $|d(s_{1},p)-d(s_{2},p)|$ and $|d(s_{1},p)-d(s_{3},p)|$ that are at most $d(s_{1},s_{2})$ and $d(s_{1},s_{3})$ respectively. Therefore, there are at most $\bigl{(}2d(s_{1},s_{2})+1\bigr{)}\cdot\bigl{(}2(d(s_{1},s_{3})+1\bigr{)}$ combinations of weights defining an additively weighted Voronoi diagram for which $p$ might be a triple intersection point, and at most two triple intersection points per diagram by Lemma 9, giving a total of at most $2\bigl{(}2d(s_{1},s_{2})+1\bigr{)}\cdot\bigl{(}2(d(s_{1},s_{3})+1\bigr{)}$ possible points that can have integer distances to all three of $s_{1}$, $s_{2}$, and $s_{3}$. ∎
The following generalization of the Erdős–Anning theorem follows immediately:
@@
If $S$ is not collinear, it has three non-collinear points to which we apply Theorem 1. ∎
@@ 3.5 Diameter bounds
We may assume without loss of generality that the unit ball of the given distance function $d$ is contained in a Euclidean unit circle. For, if it does not, let $\xi$ be the maximum Euclidean distance from the origin to any point of the unit $d$-ball, and scale both $S$ and $d$ by $1/\xi$, leaving distances between the scaled points unchanged and allowing the $d$-ball to fit within a Euclidean unit circle. Let $\rho$ be the minimum Euclidean distance from the origin of the unit circle of $d$; note that, if we treat the distance function as fixed, then $\rho$ is a constant. Then the function measured by $d$ is lower bounded by the Euclidean distance, and upper bounded by $\tfrac{1}{\rho}$ times the Euclidean distance. From the lower bound on $d$, the Euclidean diameter of $S$ is at most its $d$-diameter, $D$.
@@
Enclose $S$ in a square bounding box of (Euclidean) side length $D$, and center this within a larger square $B$ of side length $3D$, so that all points of $S$ are at Euclidean distance at least $D$ from the boundary of $B$. Construct the (Euclidean) Voronoi diagram of $S$, and let $s_{1}$ be the point of $S$ whose Voronoi cell $V_{1}$, intersected with $B$, has the smallest Euclidean area, $\leq 9D^{2}/n=O(D^{2}/n)$. Let $s_{2}$ be the Euclidean nearest neighbor of $s_{1}$ in $S$, let $m$ be the midpoint of $s_{1}$ and $s_{2}$, and construct a right triangle $ps_{1}m$ with right angle at $s_{1}$, contained in $V_{1}\cap B$, with the largest possible Euclidean area. (See Fig. 9 for an example.) We distinguish two cases:
  • If $p$ is on the boundary of $B$, as shown in Fig. 9, it is at Euclidean distance $\geq D$ from $s_{1}$. In order for right triangle $ps_{1}s_{2}$ to have area $O(D^{2}/n)$, at most the area of the set $V_{1}\cap B$ that contains it, edge $s_{1}s_{2}$ must have Euclidean length $O(D/n)$, and therefore also $d(s_{1},s_{2})=O(D/n)$ (larger than the Euclidean length by at most the constant factor $\tfrac{1}{\rho}$). For this distance to be a nonzero integer, $n$ must be $O(D)$.
  • Otherwise, $p$ is on the boundary of $V_{1}$, and equidistant (in Euclidean distance) from $s_{1}$ and some site $s_{3}$ which cannot be collinear with $s_{1}s_{2}$. Therefore, the Euclidean distance from $s_{1}$ to $p$ is at least half the Euclidean distance from $s_{1}$ to $s_{3}$. Combining this inequality with the area of the triangle and the $\tfrac{1}{\rho}$ factor by which $d$ can increase over Euclidean distance gives the bound $d(s_{1},s_{2})\cdot d(s_{1},s_{3})=O(D^{2}/n)$. It follows from Theorem 1 that $n=O(D^{2}/n)$ and therefore that $n=O(D)$.∎
@@ 4 Riemannian manifolds
One complication in these spaces is that their usual analogues of straight lines, geodesics, are only locally shortest curves between their points: two or more points may belong to a common geodesic that does not contain the shortest curve between them. Indeed, in some cases a geodesic may cross itself or form a dense subset of its surface. For this reason, we will generally work with shortest curves (between some two points of the space) rather than geodesics. These curves may not be unique: for instance, for antipodal points on a sphere, there are infinitely many shortest curves, covering the whole sphere. By the Hopf–Rinow theorem [16], every complete Riemannian manifold has a shortest curve between every two points. Thus, these are geodesic metric spaces, and in particular the Lipschitz property of distances (1) applies to them.
@@ 4.1 Examples
For any $n$, one may form a complete Riemannian manifold from $n$ spheres, connected by cutting out disks from pairs of spheres, attaching a cylindrical tube between the boundaries of the two disks, and smoothing the parts of the surface where the disks meet the tubes. This construction may be performed abstractly as a topological space; it is not necessary to embed the spheres and tubes into three-dimensional space. Let $S$ consist of one point per sphere, for spheres small enough that all distances within one sphere are less than $\tfrac{1}{3}$; then by adjusting the lengths of all the tubes, all distances between points of $S$ can be made equal to one. This construction produces arbitrarily large finite sets of points at unit distance from each other, with genus $O(n^{2})$. Because they all have the same distance, no three of these points belong to a shortest curve.
The connected sum of an infinite sequence of examples of this type for $n=4,5,6,\dots$, using tubes to connect the example for each $n$ to the examples for $n-1$ and $n+1$, forms a single complete Riemannian manifold of infinite genus in which one can find arbitrarily large (but nevertheless finite) sets of points that are all at unit distance from each other.
@@
Our proof of the Erdős–Anning theorem for strictly convex distance functions relied on the non-planarity of $K_{3,3}$. For complete Riemannian manifolds of bounded genus we need to reach further for graphs that cannot be embedded, as the following example shows.
@@
Form a complete Riemannian manifold (the hexagonal torus) by gluing opposite edges of a Euclidean regular hexagon. The hexagon’s six vertices become two points of the torus (red and blue in Fig. 11). Let $S$ consist of these two points and the center of the hexagon (yellow). Six points (gray) are equidistant from these three points. Connecting the three points of $S$ by shortest curves to these six points embeds $K_{3,6}$ on the torus.